TITLE

Thomas D. Schneider^{∗}

version = 1.00 of NAME.tex DATE

Example reference citations: [?, ?, ?]. Claude Shannon wrote this equation [?]:

H = −∑_{i} p_{i}log_{2}p_{i} (bits per symbol)  (1)

Example figure reference: Fig. ??. ⇐Fig ??

∗National Cancer Institute at Frederick, Laboratory of Experimental and Computational Biology, P. O. Box B, Frederick, MD 21702-1201. (301) 846-5581 (-5532 for messages), fax: (301) 846-5598, email: toms@ncifcrf.gov. http://www.lecb.ncifcrf.gov/˜ toms/

Figure 1: A caption.

Some example text of the figure.

REMOVE